# radarODE: An ODE-Embedded Deep Learning Model for Contactless ECG Reconstruction from Millimeter-Wave Radar

### v1

@@ IV-A Dataset and Benchmark Introduction
The dataset provided in [35] is collected by TI AWR-1843 radar with $77$ GHz start frequency and $3.8$ GHz bandwidth, providing good SNR and resolution to detect subtle vibrations that fit the proposed signal model and framework. Half of the actual dataset is released with $91$ trials for $11$ subjects, and each trial contains $3$ minutes of data (radar measurements and ECG ground truth) with $200$ Hz sampling rate collected under $4$ physiological statuses (i.e., normal breath (NB, $43$ trials), irregular breath (IB, $18$ trials), sleep (SP, $18$ trials) and post exercise (PE, $12$ trials)). In addition, the work in [35] has pre-processed the radar signal using several techniques, such as differentiator, dynamic time wrapping and clustering, to remove respiration noise preliminarily. Lastly, no existing study is found for ECG reconstruction in literature based on the same dataset, and the proposed framework MMECG in [35] will be used as the only benchmark to make a comparison with our radarODE.
In addition, the framework MMECG proposed in the same paper [35] is reproduced as the benchmark with $50$ synchronous radar signal segments as input, because MMECG cannot handle SST spectrograms. In detail, MMECG utilizes Conv1d and Transformer [60] as the decoder to simultaneously extract temporal and spatial features, and the features are further fused by multiplication and then decoded by Transconv1d and TCN.
IV-B Implementation Details of radarODE
@@
The proposed radarODE network is coded using PyTorch and trained for $200$ epochs with batch size $32$ on the NVIDIA RTX A4000 ($16$ GB) using stochastic gradient descent optimizer with early stop function [61] and learning rate $0.001$ based on a cosine annealing schedule [62]. The dataset is split into training and testing sets based on $11$-fold cross-validation for $11$ subjects to ensure the mutually exclusive between training and testing. In addition, all the ground truth characteristic peaks, PPI, and cardiac cycles are obtained by the NeuroKit2 from ECG signals [48]. Furthermore, the $4$-seconds-long input SST segments only contain the frequency component within $[1,25]$ Hz and are down-sampled to $30$ Hz in the time-axis for saving memory usage. The input raw radar signal also needs to be segmented into $4$-seconds-long pieces with a step of $0.2$ seconds, and the output ECG pieces need to be further concatenated as $3$ minutes long-term result as suggested in [35]. This research has been approved by University Ethics Committee of Xi’anJiaotong-Liverpool University with proposal number ER-SAT-0010000090020220906151929.
Figure 8: PPI estimation error based on KDE and mean value: (a) View of different subjects; (b) View of different physiological statuses; (c) CDF for the overall PPI estimation error.
IV-C Experimental Results and Evaluations
IV-C1 PPI Estimation
@@
The PPI estimation is the first module of radarODE, and the accuracy of the estimated PPI directly affects the fidelity of the concatenated morphological reference. Therefore, Figure 8a shows the PPI estimation error obtained by Algorithm 1 in terms of different subjects, using KDE or directly averaging the PPI from $50$ radar signals. The large PPI error (e.g., for subject $13$, $17$) is normally caused by the body movements or residual respiration noise due to IB or PE statuses, as also shown in Figure 8b with large median error and variation for both methods. In contrast, the subjects in NB and SP statuses tend to be stable with less body movement, and the respiration noise can be well eliminated, achieving low median PPI error as $0.03$ s and $0.02$ s using the KDE-based method for each status. Overall, it is clear that the KDE-based PPI estimation is more accurate than the mean-based estimation for each subject, as shown in the cumulative distribution function (CDF) in Figure 8c, with the overall median PPI error of $0.03$ s and $0.06$ s respectively.
Figure 9: Performance of SCEG: (a) Visualization of the rigid reconstruction (ODE Recon.) from ODE decoder with ground truth (GT) (b) Training loss comparison with benchmark and ablation study (without resampling); (c) and (d) High- and Low-fidelity of the generated morphological references (Ref.) with ground truth (GT), RMSE and correlation coefficient provided.
IV-C2 Single Cycle ECG Generator (SCEG)
The performance of the SCEG is only evaluated on the generated single cycle ECG pieces (without resampling and concatenation) in terms of root mean square error (RMSE) and Pearson-correlation coefficient (PCC), because the aim of the SCEG module is only to generate ECG pieces while the PPI estimation error in resampling stage may affect the analysis of the actual performance of SCEG (e.g., a little deviation of the peak contributes to large RMSE). The detailed experimental results and evaluations are listed as:
  1. a.
    Comparison with Benchmark and Ablation Study: The overall performance for MMECG and radarODE in single cycle ECG generation are shown in Table III, with the median RMSE/PCC of $0.091$ mV/$87.9\%$ and $0.077$ mV/$92.6\%$, respectively. In addition, to further evaluate the contributions of temporal and ODE decoders, the ablation study is implemented by only reserving one decoder without feature fusion. The results show that both decoders can work alone and achieve reasonable results as shown in Table III, but the ODE decoder has lower accuracy because it neglects the subtle ECG feature while only focusing on the characteristic peaks with rigid connections elsewhere, as shown in Figure 9a. However, the value of the ODE decoder will be revealed later in accelerating the training process and resisting body movements.
    TABLE III: Comparison with benchmark and Ablation study about temporal decoder and ODE decoder (without resampling)
    Framework Backbone Encoder Decoder Fusion Method RMSE (mV) Correlation
    MMECG [35] - Conv1d + Transformer Transconv1d + TCN Multiplication $0.091$ $87.9\%$
    SCEG DeformConv2d Conv1d Initial + Temporal - $0.086$ $89.4\%$
    Initial + ODE - $0.092$ $85.5\%$
    Initial + Temporal + ODE Multiplication + Stack $\mathbf{0.077}$ $\mathbf{92.6\%}$
    Figure 10: Corrupt ECG reconstruction: (a) and (b) Ideal reconstruction results for radarODE and MMECG for the same period; (c) Corrupt ECG reconstruction yield by MMECG with corrupt R peaks detected, and faithful reconstruction from radarODE; (c) Raw radar signal during body movements, causing the previous corrupt reconstruction; (d) Overall missed detection rate (MDR); (e) MDR under different physical statuses for radarODE and MMECG.
  2. b.
    Visualization of Training Process: Figure 9b illustrates the training loss for benchmark and ablation study with the dotted line representing the early stop of the training process and the shaded area indicating the repetition of the training process for five times. It is obvious that the ODE decoder itself cannot provide a very accurate reconstruction but could assist the entire model to achieve the lowest RMSE loss and also accelerate the convergence to achieve the RMSE of $0.17$ mV after the first epoch.
  3. c.
    Performance after Resampling and Concatenation: Although the SCEG could generate the ECG pieces with high fidelity, the resampled results will be affected by the PPI estimation error as shown in Figure 9c and 9d, because the slight deviation of the peak ruin the overall RMSE/PCC, hence requiring long-term reconstruction module to refine the concatenated ECG pieces in next step.
IV-C3 Performance of radarODE
The long-term ECG reconstruction module finally generates 3-minute-long ECG signals which will be used to evaluate the performance of the entire radarODE framework in terms of corrupt ECG reconstruction, morphological accuracy, and fine-grained cardiac feature accuracy as:
  1. a.
    Corrupt ECG Reconstruction: The ideal reconstructed ECG signals are shown in Figure 10a and 10b with corresponding RMSE/PCC labeled. However, the long-term ECG signal may contain the corrupt parts due to the presence of the body movements (especially in IB and PE), because the scale of these body movements is much more extensive than the cardiac vibrations, hence drowning the subtle cardiac features. Figure 10c shows a piece of corrupted ECG reconstruction yield by MMECG with the falsely detected R peaks noted as red dots, and Figure 10d is the corresponding raw radar signal with extensive fluctuation induced by body movements. In contrast, the radarODE could still provide faithful ECG reconstructions under body movements due to the introduction of prior knowledge (ODE model) about ECG as the morphological reference, hence gaining certain robustness in resisting noises.
    Figure 11: Morphological Accuracy for radarODE and MMECG: (a) and (b) Overall CDF of RMSE and PCC for all trials; (c) and (d) RMSE and PCC under different physical statuses; (e) and (f) RMSE and PCC across all subjects.
    In the aims of quantifying the percentage of the corrupt reconstruction, three rules are made to reject the falsely detected R peaks yielded by Neurokit2 based on the empirical observations from all kinds of corrupt samples:
@@ IV-C3 Performance of radarODE
    • The deviation of the corrupt R peak from the ground truth exceeds the absolute tolerance of $0.15$ s [35].
    • The corrupt R peak has one or several adjacent R peaks with the PPI smaller $0.3$ s.
@@
    After filtering the corrupt R peaks from all detection, the overall missed detection rate (MDR) can be calculated by $MDR=\frac{\text{Remained R Peaks}}{\text{Ground Truth R Peaks}}$ as shown in Figure 10e, with the median MDR of $0\%$ and $4.19\%$, and $90$-percentile MDR as $1.34\%$ and $4.76\%$ for radarODE and MMECG respectively, and the overall improvement of radarODE compared with MMECG is $9\%$. In addition, the CDF plots for different physical statuses are plotted as in Figure 10f with the $90$-percentile MDR of $0.12\%$, $0.85\%$, $0.32\%$, $3.71\%$ during NB, IB, SP, PE for radarODE and $3.36\%$, $6.22\%$, $0.83\%$, $9.64\%$ for MMECG. The result shows that different physical statuses have a noticeable impact on the quality of the reconstructed long-term ECG, and radarODE could provide a higher MDR than MMECG in all statuses owing to the prior knowledge in the ODE decoder.
  2. b.
@@
    Morphological Accuracy: The morphological accuracy can reveal the similarity between reconstructed and ground truth ECG signals by calculating RMSE and PCC, with RMSE mainly affected by the deviation of peaks and PCC focusing on the general shape. Figure 11a and 11b shows the overall performance of radarODE and MMECG in CDF with the median RMSE/PCC of $0.097$mV$/89.6\%$ and $0.120$mV$/81.2\%$, and $90$-percentile RMSE/PCC of $0.148$mV$/76.2\%$ and $0.213$mV$/66.3\%$, respectively. The accuracy of the long-term ECG reconstruction is obviously lower than that of the single cycle ECG pieces in Table III, because the SCEG module only needs to learn the representations in one cardiac cycle with similar ECG patterns, but the long-term ECG reconstruction module deals with signals in arbitrary time period. Another critical remark is that the overall improvement of radarODE compared to MMECG is $16\%$ and $19\%$ for RMSE and PCC across $91$ trials, indicating that the morphological prior is more helpful in generalizing the typical ECG pattern than calibrating the peaks.
@@
    In addition, Figure 11c and 11d shows the RMSE/PCC across all subjects, with radarODE always achieving better results than MMECG. In addition, the results of the long-term reconstruction show certain consistency with the previous PPI estimation error because the PPI error directly determines the fidelity of the morphological reference inputted in the long-term reconstruction module. For example, subjects $10$, $13$, $14$, and $30$ get worse results than others in either RMSE or PCC evaluation due to the large PPI estimation error as shown in Figure 8a.
@@
    Lastly, Figure 11e and 11f illustrate the RMSE/PCC for all trials in terms of different physical statuses, and radarODE outperforms MMECG across all the statuses. Additionally, the box plots for ECG recovery coincide with the PPI estimation error and show that the stable statuses (i.e., NB, SP) guarantee the reconstruction with small variance, while the unstable statues (i.e., IB, PE) may heavily ruin the radar signal in some trials due to the body movements, causing an inconsistent quality of the reconstructed ECG. In addition, the trials for PE status are the least (only $12$ trials) compared with the other three statutes, and the data shortage may also cause bad performance, because the training samples are not enough for deep neural networks to learn the pattern under PE status. However, radarODE could still provide the reconstructions with a smaller variance than MMECG, especially for unstable statues because of the morphological prior embedded in the ODE decoder.
  3. c.

### v2

@@ IV-C Dataset Description
Half of the actual dataset is released with $91$ trials for $11$ subjects (with subject ID 1, 2, 5, 9, 10, 13, 14, 16, 17, 29, 30), and each trial contains $3$ minutes of data (radar measurements and ECG ground truth) with $200$ Hz sampling rate collected under $4$ physiological statuses (i.e., normal breath (NB, $43$ trials), irregular breath (IB, $18$ trials), sleep (SP, $18$ trials) and post exercise (PE, $12$ trials)). In addition, the work in [28] has pre-processed the radar signal using several techniques, such as 3D beamforming, dynamic time wrapping, to remove respiration noise and enhance cardiac activities. Lastly, no existing study is found for ECG reconstruction in literature based on the same dataset, and the proposed framework MMECG in [28] will be used as the only benchmark to make a comparison with our radarODE.
IV-D Implementation Details and Compared Frameworks
@@
The proposed radarODE network is coded using PyTorch and trained for $200$ epochs with batch size $32$ on the NVIDIA RTX A4000 ($16$ GB) using stochastic gradient descent optimizer with early stop function [48] and learning rate $0.001$ based on a cosine annealing schedule [49]. The dataset is split into training and testing sets based on $11$-fold cross-validation with $1$ fixed subject for testing and the other $10$ subjects alternatively selected for training or validation, ensuring to make the most of all the trials while excluding the testing data from the training phase. In addition, all the ground truth characteristic peaks, PPI, and cardiac cycles are obtained by the NeuroKit2 from ECG signals [39]. Furthermore, the $4$-seconds-long input SST segments only contain the frequency component within $[1,25]$ Hz and are down-sampled to $30$ Hz in the time-axis for saving memory usage in backbone design. This research has been approved by University Ethics Committee of Xi’anJiaotong-Liverpool University with proposal number ER-SAT-0010000090020220906151929.
In addition, three frameworks are selected for comparison with the following brief introduction of the architecture:
  • MMECG [28] receives multiple 1D radar signals as input and utilizes Conv1d and Transformer as decoder to simultaneously extract temporal and spatial features. The encoded features are further fused by multiplication and then decoded by Transconv1d and TCN to produce ECG recovery.
  • RSSRNet [29] takes spectrogram (STFT) as input with a Conv2d backbone and Transformer encoder. The adopted decoder is Transconv2d, and the output is still a spectrogram and needs to be converted to the ECG signal via inverse STFT.
  • RadarNet [30] adopts 1D radar signal as input and directly generates coarse ECG signal using Conv1d. Then, several layers of ResNet are adopted to refine the ECG waveform.
V Experimental Results and Evaluations
This section provides the experimental results and evaluations in terms of three core modules as depicted in Figure 2, with the first module providing PPI estimation for input/output slicing and reshaping, the second module robustly generating ECG pieces for the single cardiac cycle, and the third module yielding the final long-term ECG recovery.
V-A Evaluations of PPI Estimation
@@
The PPI estimation is the first module of radarODE, and the accuracy of the estimated PPI directly affects the fidelity of the concatenated morphological reference. Therefore, Figure LABEL:sub@fig:ppi_obj shows the PPI estimation error obtained by Algorithm 1 for all subjects based on KDE defined in (18) or directly averaging the candidate PPI values. The large PPI error (e.g., for subject $13$, $17$) is normally caused by body movements or residual respiration noise due to IB or PE status, as also shown in Figure LABEL:sub@fig:ppi_phy with large median error and variation for both methods. In contrast, the subjects in NB and SP statuses tend to be stable with less body movement, and the respiration noise can be well eliminated, achieving low median PPI errors as $0.03$s and $0.02$s using the KDE-based method for each status.
Overall, it is clear that the KDE-based PPI estimation is more accurate than the mean-based estimation for each subject, as shown in the cumulative distribution function (CDF) in Figure LABEL:sub@fig:ppi_cdf, because the KDE-based method is robust to the outliers caused by noises and could figure out the correct PPI estimation near the majority of candidate values.
Figure 9: PPI estimation error of KDE-based and mean-based method: (a) View of different subjects; (b) View of different physiological statuses; (c) CDF for the overall PPI estimation error.
Figure 10: Performance comparison for single-cycle ECG recovery: (a) - (d) CDF of RMSE, PCC, MDR and R-peak error for all trials.
V-B Evaluations of SCEG Module in radarODE
SCEG is the core module that realizes the domain transformation and ensures the robust long-term ECG recovery in the next stage, and the performance is evaluated on the generated single-cycle ECG pieces in terms of morphological accuracy, corrupt ECG reconstruction and absolute R-peak error as shown in Figure 10 and Table VI.
V-B1 Comparison of Morphological Accuracy
The morphological accuracy is shown in Table VI as the median value of root mean square error (RMSE) and Pearson-correlation coefficient (PCC), with RMSE sensitive to the deviation of peaks and PCC focusing on the general shape. The overall performances for all trials are shown as CDF plots in Figure LABEL:sub@fig:rmse_cdf_compare and LABEL:sub@fig:sceg_cor_cdf. The results indicate that radarODE could generate high-fidelity ECG signals with good RMSE and PCC owing to the prior knowledge provided by the ODE decoder, while MMECG achieves the second-best result and shows a domain transformation ability for the majority of radar inputs with good SNR. In contrast, RSSRNet and RadarNet achieve a similar performance because they only accept single-channel inputs and cannot effectively leverage the features of $50$ channels provided in the dataset.
V-B2 Comparison of Corrupt ECG Recovery
The ability to recover ECG pieces from corrupt radar signals reveals the noise robustness of each framework, and missed detection rate (MDR) is adopted to count failed recoveries without showing characteristic ECG patterns (R peaks). Three rules are made to define the failed ECG recoveries:
@@ V-B2 Comparison of Corrupt ECG Recovery
  • The deviation of the corrupt R peak from the ground truth exceeds the absolute tolerance of $0.15$[28].
  • The corrupt R peak has one or more neighboring R peaks closer than $0.3$s.
@@
The results of MDR are shown in Table VI with CDF plots of all trials shown in Figure LABEL:sub@fig:MDR_cdf_overall. It is evident that radarODE achieves better noise robustness compared with previous work, owing to the constraint brought by ODE decoder. MMECG still has a better performance compared with RSSRNet and RadarNet due to the powerful backbone to extract features from undistorted channels. The overall performance of MDR in Figure LABEL:sub@fig:MDR_cdf_overall coincides with the morphological accuracy in Figure LABEL:sub@fig:rmse_cdf_compare and LABEL:sub@fig:sceg_cor_cdf, because the corrupt ECG recoveries also significantly affect the RMSE and PCC.
V-B3 Comparison of R-peak Timing Error
After filtering the corrupt ECG pieces, the absolute timing error of R peaks is calculated to evaluate the quality of fine-grained ECG features, with the median value and CDF plots shown in Table VI and Figure LABEL:sub@fig:sceg_r_cdf respectively. The results are still similar to previous evaluations, with radarODE recovering the most accurate R peaks and the other three frameworks showing larger R-peak deviation. In addition to the benefits brought by ODE decoder, the adopted SST inputs also provide necessary time-frequency features to help the deep learning model identify indistinctive vibrations under low-SNR scenarios.
TABLE VI: Performance Comparison for Single-cycle ECG Recovery and Ablation Study
Framework Backbone Encoder Decoder RMSE (mV) PCC MDR R Error (sec)
MMECG [28] - Conv1d + Transformer Transconv1d + TCN $0.091$ $87.9\%$ $1.24\%$ $0.012$
RSSRNet [29] Conv2d Transformer Transconv2d $0.100$ $86.0\%$ $2.17\%$ $0.019$
RadarNet [30] - Conv1d Conv1d (ResNet) $0.113$ $80.2\%$ $2.58\%$ $0.024$
radarODE Deform Conv2d Conv1d Initial + Temporal $0.086$ $89.4\%$ $1.53\%$ $0.012$
Initial + ODE $0.092$ $85.5\%$ $\mathbf{0.14\%}$ $\mathbf{0.005}$
Initial + Temporal + ODE $\mathbf{0.077}$ $\mathbf{92.6\%}$ $0.18\%$ $0.006$
Figure 11: Ablation study and visualization: (a) Rigid reconstruction (ODE Recon.) from ODE decoder with ground truth (GT); (b) Training loss comparison; (c) and (d) High- and Low-fidelity morphological references (Ref.) caused by PPI estimation error.
V-B4 Ablation Study
The ablation study is performed to further evaluate the contributions of temporal and ODE decoders. The results in Table VI reveal that both decoders can work individually and achieve reasonable results, but the ODE decoder has lower accuracy because it neglects the subtle ECG feature while only focusing on the characteristic peaks with rigid connections elsewhere, as shown in Figure LABEL:sub@fig:ode_sceg. On the contrary, the introduced ODE model could resist strong noise and achieve the lowest MDR as $0.14\%$, while the temporal decoder only gets a similar MDR ($1.53\%$) with MMECG ($1.24\%$). In radarODE, the outputs of temporal and ODE decoder are fused together to achieve noise robustness with MDR$=0.18\%$, while maintaining a faithful ECG shape (i.e., good RMSE and PCC).
Figure 12: Corrupt ECG reconstruction: (a) and (b) Ideal reconstruction results for radarODE and MMECG; (c) Corrupt ECG reconstruction yield by MMECG and faithful reconstruction from radarODE; (d) Corresponding radar signal during body movements; (e) Overall missed detection rate (MDR); (f) MDR under different physical statuses.
V-B5 Visualization of Training Process
Figure LABEL:sub@fig:loss_compare illustrates the training loss for benchmark and ablation study with the dotted line representing the early stop of the training process and the shaded area indicating the repetition of the training process for five times. From the ablation study, SCEG with only ODE decoder cannot provide a very accurate reconstruction due to the rigid shape as shown in Figure LABEL:sub@fig:ode_sceg, while the temporal decode could achieve the second-best result. After combining ODE and temporal decoders, the outputs from ODE decoder act as the morphological-prior to accelerate the convergence, achieving the RMSE of $0.17$mV after the first epoch. In addition, the morphological reference will not be destroyed by noises and could stabilize the training process, contributing to the lowest training loss as shown in Figure LABEL:sub@fig:loss_compare.
V-B6 Summary of SCEG Evaluation
The proposed SCEG module in radarODE achieves better accuracy in generating single-cycle ECG pieces with the improvements brought by the ODE decoder and SST inputs, enabling successful recoveries even under abrupt noises and achieving the best RMSE, PCC and R-peak accuracy compared with previous frameworks.
The generated ECG pieces can be resized and concatenated based on PPI estimation, but the PPI estimation error may accumulate and degrade the accuracy as shown in Figure LABEL:sub@fig:sceg_good and LABEL:sub@fig:sceg_bad, because slight deviations of the peaks ruin the overall RMSE/PCC, hence requiring long-term reconstruction module to refine the concatenated ECG pieces in the next step. To provide a comprehensive evaluation for long-term ECG recovery from various dimensions (e.g., trials, subjects, and physical statuses), MMECG is selected as the only benchmark in the next section for a clear result visualization.
V-C Overall Evaluations of radarODE
The long-term ECG reconstruction module finally generates 3-minute-long ECG signals for the evaluations of the entire radarODE framework in terms of corrupt ECG reconstruction, morphological accuracy, and fine-grained cardiac feature accuracy.
V-C1 Corrupt ECG Reconstruction
The ideal reconstructed ECG signals are shown in Figure LABEL:sub@fig:overall_ode and LABEL:sub@fig:overall_mmecg with corresponding RMSE/PCC labeled. However, the long-term ECG signal may contain corrupt parts due to the presence of body movements (especially in IB and PE). Figure LABEL:sub@fig:overall_noise shows the corrupted ECG reconstruction yield by MMECG with the falsely detected R peaks noted as red dots, and Figure LABEL:sub@fig:overall_rcg is the raw radar signal with extensive distortion induced by body movements. In contrast, the radarODE could still provide faithful ECG reconstructions under body movements due to the introduction of prior knowledge (ODE model) about ECG as the morphological reference, hence gaining certain robustness in resisting noises.
Figure 13: Morphological accuracy comparison: (a) and (b) Overall CDF of RMSE and PCC for all trials; (c) and (d) RMSE and PCC under different physical statuses; (e) and (f) RMSE and PCC across all subjects.
@@ V-C1 Corrupt ECG Reconstruction
The overall MDR can be calculated from the recovered ECG signal with the CDF plot shown in Figure LABEL:sub@fig:detection_cdf_overall, and the overall improvement achieved by radarODE is $9\%$. In addition, the CDF plots for different physical statuses are plotted in Figure LABEL:sub@fig:detection_cdf_phy with the $90$-percentile MDR of $0.12\%$, $0.85\%$, $0.32\%$, $3.71\%$ during NB, IB, SP, PE for radarODE and $3.36\%$, $6.22\%$, $0.83\%$, $9.64\%$ for MMECG. The result shows that different physical statuses have noticeable impacts on the quality of the reconstructed long-term ECG, and radarODE could provide a lower MDR than MMECG in all statuses owing to the prior knowledge in the ODE decoder.
V-C2 Morphological Accuracy
@@
The morphological accuracy measures the similarity between reconstructed and ground truth ECG signals by calculating RMSE and PCC. Figure LABEL:sub@fig:rmse_cdf and LABEL:sub@fig:cor_cdf show the overall performance of radarODE and MMECG in CDF with the median RMSE/PCC of $0.097$mV$/89.6\%$ and $0.120$mV$/81.2\%$. It is worth noticing that the overall improvement of radarODE compared to MMECG is $16\%$ and $19\%$ for RMSE and PCC across $91$ trials respectively, indicating that the morphological-prior is more helpful in generalizing the typical ECG pattern than calibrating the peaks.
@@
In addition, Figure LABEL:sub@fig:rmse_obj and LABEL:sub@fig:cor_obj illustrate the RMSE/PCC across all subjects, with radarODE always achieving better results than MMECG. It is worth noting that the results of the long-term reconstruction show certain consistency with the previous PPI estimation error, because the fidelity of the morphological reference is directly affected by the PPI error. For example, subjects $10$, $13$, $14$, and $30$ get worse results than others in either RMSE or PCC evaluation due to the large PPI estimation error as shown in Figure LABEL:sub@fig:ppi_obj.
@@
Lastly, Figure LABEL:sub@fig:rmse_phy and LABEL:sub@fig:cor_phy illustrate the RMSE/PCC for all trials in terms of different physical statuses, and the box plots show that the stable statuses (i.e., NB, SP) guarantee the reconstruction with small variance. In contrast, unstable statues (i.e., IB, PE) can severely ruin the radar signal due to body movements, causing an inconsistent quality of the reconstructed ECG. However, radarODE could still provide the reconstructions with a smaller variance than MMECG, especially for unstable statues because of the morphological-prior embedded in the ODE decoder.
V-C3 Fine-Grained Cardiac Events Reconstruction